OpenStudy (anonymous):

Graph the system of quadratic inequalities. y leq −x2 + 3 y > x2 − 4x − 2 Which x−value is part of the solution?

OpenStudy (anonymous):

\[y \le -x^2+3,~or~x^2\le -y+3 ...(1)\] it is a downward parabola. vertex is (0,3), cuts x-axis at \[-\sqrt{3}~and~ \sqrt{3}\] put x=0 and y=0 in (1) \[0\le0+3,~or~0\le3\]which is true. Hence origin lies in the graph. \[y>x^2-4x-2\] \[or~y>x^2-4x+(\frac{ 4 }{ 2 })^2-\left( \frac{ 4 }{2 } \right)^2-2\] \[or~y>\left( x-2 \right)^2-4-2\] \[or~y+6>\left( x-2 \right)^2 ....(2)\] \[or~\left( x-2 \right)^2<y+6\] it is an upward parabola with vertex at(2,-6) it cuts x-axis when y=0 \[\left| x-2 \right|<\sqrt{6},\] \[-\sqrt{6}<x-2<\sqrt{6} ,~or~2-\sqrt{6}<x<2+\sqrt{6}\] put x=0,y=0 in (2) 6>4 which is true. Hence origin lies in the graph.  shade portion is the required region.